s = ?

s = ?


 

Grade:11

3 Answers

samantha k
12 years ago

s=ut+1/2(at2)

nihanth g
12 years ago

s=ut+1/2at2

SHUBHAM PAUL
12 years ago

WHEN ACCELERATION OF OBJECT IS CONSTANT ,THEN WE CAN WRITE ,S=UT+1/2AT2 WHERE 'S' IS THE DISPLACEMENT OF BODY AT TIME [T] .SO YOU HAVE TO MENTION THE WHOLE CONDITION OF MOTION.